# Effective and Transparent RAG: Adaptive-Reward Reinforcement Learning for Decision Traceability

### v1

@@ Appendix C Details of Evaluation Metrics
C.1 Exact match and normalization
@@ C.3 LLM-as-a-Judge evaluation
Appendix D Details of Supervised Fine-Tuning
@@
Appendix E Details of Training Configuration
We train all models using the Open-R1 reinforcement learning framework with DeepSpeed ZeRO Stage 2 optimization and bfloat16 mixed-precision. The total batch size is 256. The learning rate is set to 3e-6. We generate 7 rollout samples per input for reward estimation. We set temperature = 0.9 during rollout, KL coefficient $\beta=0.04$, number of iterations per batch $\mu=1$, and clipping parameter $\epsilon=0.2$.
Appendix F KL Divergence Stabilization
@@
Appendix G Additional Experiemntal Results on Reward Variants
To further evaluate the impact of different accuracy reward designs during training, we conduct the same experiments on 2WikiMultiHopQA and MuSiQue. Results are shown in Appendix G and Appendix G. On 2WikiMultiHopQA, EM-based reward continues to yield the best performance across all three metrics, confirming its generalizability on structured reasoning tasks.
On MuSiQue, however, we observe that GPT- and F1-based judgment outperform EM-based reward. This may be attributed to the greater difficulty of the MuSiQue dataset—each question is paired with 20 retrieved paragraphs instead of 10—making exact matching harder and rendering fuzzy evaluation criteria more helpful for guiding learning. These findings suggest that while EM-based reward is generally more stable and interpretable, looser metrics might be preferable in high-noise or high-ambiguity retrieval settings.
Table 10: Reward method comparison on 2WikiMultiHopQA. EM-based reward remains the most effective.
Judge Method EM F1 LJ
GPT-4o 0.644 0.730 0.740
F1 0.646 0.736 0.732
EM (ours) 0.660 0.752 0.774
Table 11: Reward method comparison on MuSiQue. F1- and GPT-based rewards show advantage in this more challenging setting.
Judge Method EM F1 LJ
GPT-4o 0.430 0.533 0.516
F1 0.438 0.548 0.502
EM (ours) 0.400 0.520 0.508

### v2

@@ Appendix D Details of Evaluation Metrics
D.1 Exact match and normalization
@@ D.3 LLM-as-a-Judge evaluation
Appendix E Details of Supervised Fine-Tuning Baselines
@@
Appendix F Details of Training Configuration
We train all models using the Open-R1 reinforcement learning framework with DeepSpeed ZeRO Stage 2 optimization and bfloat16 mixed-precision. The total batch size is 256, learning rate is set to 3e-6. We generate 7 rollout samples per input for reward estimation. We set temperature = 0.9 during rollout, KL coefficient $\beta=0.04$, number of iterations per batch $\mu=1$, and clipping parameter $\epsilon=0.2$. Our experiments were conducted using 8×A100-80G GPUs for approximately 20 hours.
Appendix G Empirical Analysis of KL Estimators
@@
Appendix H Cross-task and Cross-scale Generalization
H.1 Small-scale Model Experiments
To explore the feasibility of our method under more resource-constrained settings, we conducted experiments using Qwen2.5-3B-Instruct, a smaller backbone model. The evaluation results on the HotpotQA dataset are shown in Table 8.
Table 8: Performance of ARENA on the 3B-scale model.
Model EM F1 Format Relevance
Qwen2.5-3B-Instruct 37.2 50.0 90.2 47.5
ARENA-3B (ours) 52.8 67.6 100.0 81.2
ARENA-3B demonstrates substantial improvements in both EM and F1 over the untrained version, while also enhancing format adherence and citation accuracy.
H.2 Task Transfer to Summarization
To further test the generality of our approach, we conducted preliminary experiments applying ARENA to a summarization task. Following our discussion, we hypothesized that task-specific, structured reward functions can generalize beyond QA.
We collected several summarization datasets from the Super-NaturalInstructions benchmark (Wang et al., 2022), including topic-focused news summarization (Narayan et al., 2018). Using GPT-4o (version 2024-11-20), we annotated relevant source sentences for each reference summary. This enabled us to construct structured triples for RL training: <relevance> <analysis> <summary>, analogous to our QA setting.
We trained models on 5,000 examples and evaluated on 500 held-out samples. Results are presented in Table 9.
Table 9: Summarization performance using structured RL with customized rewards.
Model R-1 R-2 R-L R-avg Relevance
Qwen2.5-3B-Instruct 22.6 4.3 15.3 14.1
ARENA-3B (RAG) 23.7 4.4 16.3 14.8 46.6
SFT-3B 25.5 5.8 17.5 16.3
Naive-GRPO-3B 28.9 7.2 20.8 19.0
ARENA-summary-3B (ours) 29.7 7.6 21.4 19.6 53.0
As shown above, ARENA-summary-3B achieves the best performance across all ROUGE metrics, confirming that our reward-guided RL framework effectively extends to summarization. Notably, even ARENA-3B trained on QA improves over the base model, indicating promising cross-task transferability.
Appendix I Ablation on the Bonus Reward
To assess the contribution of the bonus reward, we conduct a focused ablation where we retain the format, accuracy, and relevance reward components, but remove the bonus reward that is normally applied when all three criteria are simultaneously satisfied.
While this ablated model (ARENA-7B-nobonus) continues to improve on accuracy metrics (EM and F1), we observe a clear degradation in structural metrics over time. As shown in Table 10, both format and relevance scores decline substantially after step 75, and the final model fails to consistently generate well-structured and attributed outputs. These results demonstrate the crucial role of the bonus reward in maintaining output consistency throughout training.
Table 10: Training trajectory of ARENA without the bonus reward (evaluated on HotpotQA).
Variant EM F1 Format Relevance
ARENA-7B-nobonus_checkpoint-25 58.4 73.2 99.4 86.2
ARENA-7B-nobonus_checkpoint-50 58.8 73.2 95.6 81.9
ARENA-7B-nobonus_checkpoint-75 58.2 73.2 91.0 29.5
ARENA-7B-nobonus (step 95) 60.4 74.5 86.0 13.6
Appendix J Comparing ARENA-7B with Advanced Closed-Source Reasoning Models
To better understand the performance gap between open-source 7B models and leading closed-source resoning models, we conduct pilot experiments across several LLMs under two inference settings: direct, where the model answers based solely on the question, and direct RAG, where the model receives both the question and retrieved context passages as input.
As shown in Table 11, models such as GPT-4o, OpenAI-o1, and DeepSeek-R1 significantly outperform open 7B models. These results reveal a substantial gap in RAG generation quality and highlight the need to improve the answer generation capabilities of smaller open-source models.
According to Table 12, our ARENA-7B achieves performance comparable to GPT-4o and DeepSeek-R1 across all evaluation metrics, and substantially narrows the gap with OpenAI-o1. This demonstrates the effectiveness of our structured RL training in improving output quality, making smaller open-source models more competitive for multi-hop reasoning.
Table 11: Performance of vanilla models under two inference strategies: direct and direct RAG. All models are evaluated using identical retrieved contexts. Bold indicates the highest value per column. Metrics: EM = Exact Match, F1 = F1 score, LJ = LLM-as-a-Judge.
Model HotpotQA 2Wiki MuSiQue
EM F1 LJ EM F1 LJ EM F1 LJ
Direct
Qwen2.5-7B-Instruct 18.8 24.7 26.2 22.2 25.3 23.4 4.0 9.0 6.1
Llama3.1-8B-Instruct 19.4 28.3 28.4 21.0 24.6 23.2 3.8 8.0 5.4
DeepSeek-R1-Distill-Qwen-7B 8.6 15.0 16.2 14.4 20.6 23.4 1.2 4.3 2.6
Qwen3-8B 24.8 34.4 36.7 22.0 26.3 26.4 3.0 9.2 6.0
GPT-4o 37.8 50.0 50.0 34.2 38.6 36.6 10.0 19.2 12.4
OpenAI-o1 51.0 66.7 72.6 53.0 63.4 68.4 29.2 40.4 39.2
DeepSeek-R1 42.2 56.4 59.0 45.4 53.6 54.6 20.2 33.0 29.4
Direct RAG
Qwen2.5-7B-Instruct 48.4 62.8 66.0 33.4 42.4 41.2 25.2 35.4 30.6
Llama3.1-8B-Instruct 52.8 67.6 73.8 39.8 47.8 48.4 24.8 37.0 32.4
DeepSeek-R1-Distill-Qwen-7B 33.2 48.7 71.2 29.0 40.7 65.8 11.6 18.4 27.8
Qwen3-8B 58.2 71.9 76.8 65.2 72.7 78.4 33.6 39.7 39.4
GPT-4o 62.8 78.8 82.0 60.6 68.7 71.7 50.5 61.7 59.3
OpenAI-o1 65.6 81.6 89.0 70.2 80.5 83.7 62.2 74.4 76.2
DeepSeek-R1 61.8 78.1 83.6 72.1 80.1 83.8 55.9 70.2 70.3
Table 12: Comparison between ARENA-7B and leading closed-source models on HotpotQA.
Model EM F1 LJ
GPT-4o 62.8 78.8 82.0
OpenAI-o1 65.6 81.6 89.0
DeepSeek-R1 61.8 78.1 83.6
ARENA-7B 62.8 76.2 81.2
Appendix K Hop-level Accuracy Analysis
We analyzed model performance across different reasoning hop levels. The results are summarized in Table 13. As expected, questions with fewer reasoning hops generally yield higher accuracy and relevance, while those with more hops—particularly in MuSiQue—pose significant challenges. These results highlight the difficulty of multi-hop retrieval and reasoning.
Table 13: Accuracy of ARENA-7B across different reasoning hop levels.
Dataset Hops Total Relevance Answer All Correct
HotpotQA 2 500 416 / 83.2% 314 / 62.8% 272 / 54.4%
2Wiki 2 406 360 / 88.7% 249 / 61.3% 232 / 57.1%
2Wiki 4 94 49 / 52.1% 81 / 86.2% 47 / 50.0%
MuSiQue 2 251 125 / 49.8% 109 / 43.4% 76 / 30.3%
MuSiQue 3 153 4 / 2.6% 55 / 36.0% 3 / 2.0%
MuSiQue 4 96 1 / 1.0% 36 / 37.5% 0 / 0.0%